OpenStudy (anonymous):

What is the simplified form of 15 times x to the fifth power over 20 times y to the sixth power times the fraction 6 times y to the fourth power over 5 times x-squared ?

OpenStudy (anonymous):

\[15x^6/20y^5 * 6y^2/5x^4\]

OpenStudy (anonymous):

@phi @tkhunny Hey guys, can yall help me with this, I've been stuck on it for like 20 minutes and my lesson really isn't of any help. Thanks

OpenStudy (tkhunny):

Well, your words don't match your notation. The exponents are switched in both fractions.

OpenStudy (anonymous):

Oh woops, well the proper one is the equation itself.

OpenStudy (anonymous):

The answer choices are: \[9x^2/10y^3\] \[9x^3/10y^2\] \[10x^3/9y^2\] \[10x^2/9y^3\]

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@amistre64 @dan815

OpenStudy (anonymous):

What is the equation written properly?

OpenStudy (anonymous):

nvm i got it

OpenStudy (anonymous):

It's the first reply, right under the original question posted.

OpenStudy (anonymous):

The exponents did not match each variable.

OpenStudy (anonymous):

What would the first step be?

OpenStudy (anonymous):

You're welcome.

OpenStudy (tkhunny):

15x^6/20y^5∗6y^2/5x^4 Rearrange \(\dfrac{15\cdot 6}{20\cdot 5}\cdot\dfrac{x^{6}}{x^{4}}\cdot\dfrac{y^{2}}{y^{6}}\) Does that look any easier?

OpenStudy (anonymous):

wouldn't consider what I said rudeness, simply looking for a straight forward answer in a timely manner. AndI'm sorry, but not much @tkhunny I got the answer from the other guy's link, but if you walk me through how you do it, it would be greatly appreciated.

OpenStudy (amistre64):

its a concept called "common factors", and something called an inverse property. not too sure where the disconnect would be without guiding you thru it with leading questions tho.

OpenStudy (amistre64):

the variable parts are the simplest .... x^k and x^n have xs in common the number parts would have to be factored of simplified in some manner to see what they have in common